1. (a) if f(t)=e^-wt sin wt, show that d^2f/dt^2 +2w df/dt + 5wf=0

(b) A curve passes through the points (1, 2) and its gradient function is 3x^4-1 all over X^2. Find the equation of the curve.


1
Expert's answer
2021-07-05T09:03:49-0400

(a)

"f'(t)=-we^{-wt}\sin wt+we^{-wt}\cos wt"

"f''(t)=w^2e^{-wt}\sin wt-w^2e^{-wt}\cos wt"

"-w^2e^{-wt}\cos wt-w^2e^{-wt}\sin wt"

"=-2w^2e^{-wt}\cos wt"

"\dfrac{d^2f}{dt^2}+2w\dfrac{df}{dt}+5wf"

"=-2w^2e^{-wt}\cos wt+2w(-we^{-wt}\sin wt+we^{-wt}\cos wt)"

"+5w(e^{-wt}\sin wt)=-2w^2e^{-wt}\sin wt+5we^{-wt}\sin wt"

"=-2w^2e^{-wt}\sin wt+5we^{-wt}\sin wt\not=0"

(b)


"f(x)=\int(3x^4-1)dx=\dfrac{3}{5}x^5-x+C"

"f(1)=\dfrac{3}{5}(1)^5-1+C=2=>C=\dfrac{12}{5}"

"f(x)=\dfrac{3}{5}x^5-x+\dfrac{12}{5}"